Question 4 0/ 1 pts Your machine operator is becoming more productive and generates additional proft each year. You expect to receive $27,000 in profit at the end of the year, but this will increase by 4% a year for the next 12 years. If you have a 16% annual interest rate on your investments (compounded annually), what is the Future Value of the profnts at the end of 12 years from now? 164314
Answer #1


Profit in year 1 (A1) = $27,000

Constant rate of increase in profit (g) = 4%

Annual interest rate (i) = 16%

Time period (n) = 12 years

STEP 1

Calculate the Present Value -

Pg = A_{1}left [ rac{1-left ( rac{1+g}{1+i} ight )^{n}}{i-g} ight ]

1- (1+00412 0.16-0.04 Pa 27000

1.04 12 7 Pa 27000 0.12

Py-20 1- 0.2697 0.12

0.7307 0.12 Pg 27000

Pg = 164,316.6

STEP 2

Calculate the Future Value -

FV = Pg(F/P, i, n)

FV = 164,316.6 (F/P, 12%, 12)

FV = 164,316.6 * 3.8960

FV = 640,177.47

So,

The Future Value of the profits is $640,177.47